A balloon is filled with 2000 mL of air at a pressure of 2 atm. What is the volume of the balloon if the pressure is changed to 5 atm
1
Expert's answer
2020-04-22T09:17:20-0400

p1V1 = p2V2

p2 = p1V1/V2 = 2*2000/5 = 800 ml